# DynaSpec: Context-aware Dynamic Speculative Sampling for Large-Vocabulary Language Models

### v2

@@ 5.1 Experimental Settings
0: prefix sequence $c$, speculation length $\gamma$, target model $\mathcal{T}$, router $r_{\theta}$, drafter $\mathcal{D}$, LM head $W_{LM}$, clusters $\{C_{m}\}$ over vocabulary, first token cluster budget $k_{\mathrm{first}}$, second token cluster budget $k_{\mathrm{second}}$, token budget $k_{\mathrm{t}}$, empty draft list $d$, empty draft list score $\mathrm{d\_scores}$
@@
22: verify once with $\mathcal{T}$ over full $\mathcal{V}$ on the top-$k$ subset of $d$;
5.2 Mean Acceptance Length
Table 1 summarizes mean acceptance lengths across seven benchmarks using Llama-3-8B. DynaSpec improves average acceptance from $3.64$ (FR-Spec) to $3.79$ tokens per verification step, about a 4.1% relative gain, while using a smaller average shortlist of roughly 27K tokens compared to FR-Spec’s fixed 32K. This indicates that context-aware routing recovers most of the full-vocabulary performance ($3.89$) at substantially lower draft-time cost. Consistent with observations from EAGLE-2, tasks such as QA and summarization exhibit shorter accepted lengths. In contrast, code generation, where rare tokens are crucial, shows the largest gap: FR-Spec drops to $3.31$, while DynaSpec achieves $3.85$, nearly matching full-vocabulary decoding ($3.92$). These results highlight that dynamic, context-dependent shortlisting preserves acceptance quality even under vocabulary reduction. Similar trends hold for Qwen-2-7B (Table 1), confirming the robustness of the method across architectures.
5.3 Decoding Speed
We evaluate end-to-end decoding throughput using the CUDA-based implementation of FR-Spec and our plug-in meta-classifier extension. Results are summarized in Table 2. DynaSpec achieves higher token-generation throughput than FR-Spec on four of the seven datasets, confirming that dynamic, context-aware shortlisting effectively reduces draft-time overhead. For the remaining tasks, slightly lower speeds are observed, which we attribute to the current implementation of the sub-sampled (indexed) matrix multiplication kernel. Further kernel-level optimization is expected to close this gap. Overall, DynaSpec demonstrates consistent efficiency gains across diverse workloads while maintaining full-vocabulary exactness during verification.
5.4 Position-Aware Vocabulary Shortlisting
For the position-aware strategy, the results are presented in Figure 3. Using a position-aware subset size outperforms a fixed size across all draft steps, as the first few tokens have a disproportionate impact on acceptance and throughput, an observation consistent with prior findings in Gumiho li2025gumiho. Allocating a larger cluster budget to early tokens preserves acceptance where it matters most, while decaying the shortlist later reduces indexed-GEMM cost without affecting verification. As shown in Figure 3, this adaptive schedule yields higher mean accepted lengths at comparable or smaller shortlist sizes. Table 3 further reports results for combining the position-aware schedule with FR-Spec, which likewise improves mean acceptance and reduces average shortlist size. Overall, these results confirm that position awareness complements context-aware routing by focusing computation on the most influential draft steps and amortizing overhead in later ones.
6 Conclusion
[…]
Appendix A Appendix
A.1 Position-aware frequency-ranked subset (PA-FR)
@@
Table 3: Mean accepted length comparison for Fixed top-$k$ tokens represented by FR-Spec-F vs. Position-aware top-$k$ tokens represented by FR-Spec-PA across seven datasets using Llama-3-8B model. Vocab size represents the mean shortlisted vocab size.
A.2 Meta classifier training
To train the meta classifier, we run eagle-2 li2024eagle2 on subsets of ShareGPT and UltraChat200K ding2023enhancing, and collect the hidden states of the draft model, token embeddings, and the top-k tokens at each draft step. Using these, we construct the training dataset for the meta-classifier as $\{(h_{i-1},embed(x_{i})),map2meta(top\_k\_tokens_{i})\}$, where $h_{i}$ is the hidden state of the draft model at step $i$, $embed(x_{i})$ is the embedding of token $x_{i}$, and $map2meta(top\_k\_tokens_{i})$ represents the cluster IDs obtained by mapping the top-k tokens to their clusters.

### v3

@@ A.1 Dynaspec algorithmic procedure
0: prefix sequence $c$, speculation length $\gamma$, target model $T$, router $r_{\theta}$, drafter $D$, LM head $W_{LM}$, clusters $\{C_{m}\}$ over vocabulary, first token cluster budget $k_{\mathrm{first}}$, second token cluster budget $k_{\mathrm{second}}$, token budget $k_{\mathrm{t}}$, empty draft list $d$, empty draft list score $\mathrm{d\_scores}$
@@
22: verify once with the target model $T$ over full $\mathcal{V}$ on the top-$k$ subset of $d$;
A.2 Proofs
Theorem A.1 (Distribution overlap lower bound).
For any context $c$,
$A(c)\geq(1-\beta)\left(p_{c}(S)-\textnormal{TV}(p_{c}(\cdot\mid S),q_{c}(\cdot\mid S))\right).$
Proof.
Starting from the overlap identity
$A(c)=\sum_{x\in V}\min(p_{c}(x),\tilde{q}_{c}(x))\geq\sum_{x\in S}\min(p_{c}(x),\tilde{q}_{c}(x)).$
Since, for $x\in S,\tilde{q}_{c}(x)\geq(1-\beta)q_{S,c}(x)$, then
$A(c)\geq\sum_{x\in S}\min(p_{c}(x),\tilde{q}_{c}(x))\geq(1-\beta)\sum_{x\in S}\min(p_{c}(x),q_{S,c}(x)).$
Writing $p_{c}(x)=p_{c}(S)p_{c}(x\mid S)$. Also, $q_{S,c}(x)=q_{c}(x\mid S)$ by definition. Then
$A(c)\geq(1-\beta)\sum_{x\in S}\min(p_{c}(S)p_{c}(x\mid S),q_{c}(x\mid S)).$
Also, for each $S$,
$\min(p_{c}(S)p_{c}(x\mid S),q_{c}(x\mid S))\geq p_{c}(S)\min(p_{c}(x\mid S),q_{c}(x\mid S)).$
Therefore, we have
$A(c)\geq(1-\beta)p_{c}(S)\sum_{x\in S}\min(p_{c}(x\mid S),q_{c}(x\mid S)).$
Using the fact that $\sum_{x\in S}\min(a_{x},b_{x})=1-\text{TV}(a,b)$ for distributions on $S$. This gives
$A(c)\geq(1-\beta)p_{c}(S)(1-\text{TV}(p_{c}(.\mid S),q_{c}(.\mid S))).$
Also, since $p_{c}(S)\leq 1$
$p_{c}(S)(1-\text{TV}(p_{c}(.\mid S),q_{c}(.\mid S))\geq p_{c}(S)-\text{TV}(p_{c}(.\mid S),q_{c}(.\mid S)),$
which completes the proof. ∎
The Table 3 below shows the empirical computation of the quantities of interest in the above lower bound and respective comparison between Fr-Spec and Dynaspec.
Table 3: Computation theoretical metrics on the HumanEval (Code) dataset, evaluated over 500 contexts. DynaSpec adopts the 28k vocabulary size.
Metric Value
$p_{c}(S_{\text{FR-Spec}})$ 0.8544
$p_{c}(S_{\text{DynaSpec}})$ 0.87435
$TV_{\text{FR\_Spec}}$ 0.88697
$TV_{\text{DynaSpec}}$ 0.87447
$p_{c}(S_{\text{FR\_Spec}})-TV_{\text{FR\_Spec}}$ 0.0325-0.0325
$p_{c}(S_{\text{DynaSpec}}-TV_{\text{DynaSpec}}$ 0.0001-0.0001
Theorem A.2 (Expected accepted length).
Consider a block proposal of length $b$, and let the contexts along the draft be $c_{1},c_{2},\ldots,c_{b}$ such that each depends on previously proposed tokens. Let $A_{i}\coloneqq A(c_{i})$ be the overlap at step $i$. Let $L\in\{0,1,\ldots,b\}$ be the number of consecutively accepted draft tokens before the first rejection. If each step uses a proposal $\tilde{q}_{c_{i}}$ and the acceptance test is the standard speculative accept/reject, so that the probability of accepting token $i$ given it was reached is atleast $A_{i}$, then
$\mathbb{E}[L]\geq\sum_{i=1}^{b}\prod_{j=1}^{i}A_{j}.$ (7)
Proof.
Let $E_{i}$ represent the event that the first $i$ draft tokens are accepted. Then $L\geq i\iff E_{i}$ occurs. Therefore,
$\mathbb{E}[L]=\sum_{i=1}^{b}Pr(L\geq i)=\sum_{i=1}^{b}\text{Pr}(E_{i}).$
By chain rule,
$\text{Pr}(E_{i})=\prod_{j=1}^{i}\text{Pr}(\text{accept }j\mid E_{j-1}).$
Under the standard speculative accept/reject criterion, a tight lower bound on $\text{Pr}(\text{accept }j\mid E_{j-1})$ is the overlap $A_{j}$. This gives the desired result. ∎
A.3 Position-aware frequency-ranked subset (PA-FR)
@@
Table 4: Mean accepted tokens comparison for Fixed top-$k$ tokens represented by FR-Spec-F vs. Position-aware top-$k$ tokens represented by FR-Spec-PA across seven datasets using Llama-3-8B model. Vocab size represents the mean shortlisted vocab size.